# FiniteTimeBibliography.bib

### v1

@@ 5.1 Completely positive time averages
    Through a relation between non-negativity conditions and Fourier transforms [FTpositivity2014, FTpositivityConvex], this implies:
@@
    $\widetilde{w}_{+}(\delta E)\geq 0,\text{ for all }\ \delta E\in\mathbb{R}.$ (83)
@@
where $\overline{\Theta}(x)=1$ if $x$ is true and 00 otherwise. However, the conventional time average $w(t)=(2T)^{-1}\overline{\Theta}(\lvert t\rvert\leq T)$ is not completely positive. Another important class of functions that are completely positive, which we will not discuss until Sec. 8, are discrete-time versions of Eq. (84) which are more accessible in an experimental setting.
Additionally, for some $10<W<$ 1, let $\Delta E_{W}>0$ be such that:
@@ 5.2 Energy-band thermalization from autocorrelators
Now, we will explicitly show that energy-band thermalization can be inferred from autocorrelators with a completely positive weighting function $w_{+}(t)$ which could, in principle, have support only on a finite range (or even finite set) of times. The intuition behind this relation has been described following Eq. (78), and we begin here by filling in the technical details. Formally, we will adapt a proof strategy that appeared in Refs. [ZelditchTransition, Sunada, Zelditch], generalize it to time averages with weights $w_{+}(t)$ and energy-band thermalization in place of (weak) eigenstate thermalization, and excise from it all references to the classical limit or limits such as $T\to\infty$ (or $\Delta E\to 0$). The following proposition results:
@@ 5.3 Bypassing energy levels: Global thermalization from autocorrelators
We emphasize, however, that it may still be possible for a complete basis of states that is not an eigenbasis of $\hat{\Pi}_{A}$ to thermalize this observable without implying the thermalization of the autocorrelator, nor therefore the thermalization of all other states. This strongly implicates $\hat{\rho}_{A}$ or any of its eigenbases as the most direct determiners of thermalization — which is a more accessible set of states than the complete set of all product states implicated in a slightly different setting [ETHNecessity] — while the observable may continue to thermalize in certain other bases of states without any implications for global thermalization [rigol2008eth, HarrowHuangWithoutETH].
@@ 6.1 Motivation: general structure of accessible thermalization
To motivate our strategy, let us consider how to constrain the matrix elements of $(\hat{\Pi}_{A}-\langle\hat{\Pi}_{A}\rangle_{\Sigma_{d}}\hat{\mathds{1}})$ within an energy-band in $\Sigma_{d}$ given its expectation value in an arbitrary “operator” $\hat{\gamma}$:
@@
Generically, $\hat{\Pi}_{A}$ has nonvanishing matrix elements between any two energy levels in $\mathcal{H}$. Further, individual off-diagonal matrix elements are typically inaccessible for both $\hat{\gamma}$ and $\hat{\Pi}_{A}$. To obtain a rigorous inequality for the subset of matrix elements within the energy band without having detailed knowledge of the other off-diagonal elements, the easiest strategy is to ensure that each term on the right hand side is non-negative, so that the subset of energy-band terms is always constrained to be no greater than the expectation value on the left hand side. This requires precise phase cancellations between the matrix elements of $\hat{\gamma}$ and $(\hat{\Pi}_{A}-\langle\hat{\Pi}_{A}\rangle_{\Sigma_{d}}\hat{\mathds{1}})$:
@@
Crucially, Eq. (99) appears to require $\hat{\gamma}$ to be constructed out of $(\hat{\Pi}_{A}-\langle\hat{\Pi}_{A}\rangle_{\Sigma_{d}}\hat{\mathds{1}})$ (to ensure that $\hat{\gamma}$ knows about the phase of the right hand side) sandwiched between suitably chosen “phase-cancelling” functions of $\hat{H}$ (so that one can implement dynamics or restrict to an energy shell without altering the phase). This appears to be a natural way to ensure this phase cancellation without any detailed analytical knowledge of $\hat{\Pi}_{A}$, beyond the ability to prepare/measure linear functionals of this observable. Given some (possibly multidimensional) parameter $s$, the preceding argument suggests the general structure
$\hat{\gamma}=\frac{1}{\operatorname{Tr}[\hat{\Pi}_{A}]}\int\mathrm{d}s\ \xi(s)\ f_{s}(\hat{H})\left(\hat{\Pi}_{A}-\langle\hat{\Pi}_{A}\rangle_{\Sigma_{d}}\hat{\mathds{1}}\right)g_{s}(\hat{H})$ (100)
@@ 6.2 Echo dynamics for accessing energy shells
As noted in Sec. 1, $L_{AA}(t_{1},t_{2})$ is a coarse-grained variant of the Loschmidt echo, which measures quantum interference effects between two possible evolution times of the state $\hat{\rho}_{A}$ as witnessed by the observable $\hat{\Pi}_{A}$. $L_{A}(t)$ is a stranger quantity that acts on the initial state $\hat{\rho}_{A}$ but has no observable, and we would like to refer to it as a “one-sided” echo; alternatively, it can be interpreted as an echo of the maximally mixed state with the observable $\hat{\Pi}_{A}$. Finally, $L_{\mathcal{H}}(t)$ is the spectral function of the Hamiltonian $\mathcal{H}$, measuring the Fourier transform of its probability density of energy levels (normalized to $1$); its squared magnitude is the spectral form factor [Haake, BerrySpectralRigidity]:
@@
Finally [^15] Here, we assume that $\mathcal{E}_{V}$ will typically be a closed set, while we have taken the energy-band to be an open set/interval; this is purely a choice of convention., given some $V$ such that $10<V<$ 1 let us introduce $\mathcal{E}_{V}$ as a set of energy values containing $E_{c}$ in which $\widetilde{v}_{+}(E-E_{c})$ is as large as $V$:
@@ 6.3 Energy shell thermalization from quantum dynamical echoes
in a system of $N$ particles, for some $\alpha>0$, for accessibility purposes. Intuitively, this is in line with what we saw for global thermalization in Sec. 5: a vanishingly small fraction of states may fail to thermalize with accessible resources (Sec. 4) even with autocorrelator thermalization, and to have higher resolution and restrict this fraction (perhaps all the way to $<1/D$) one must allow “inaccessible” values of $\epsilon_{A}$. Here, we see that the energy shell $\Sigma_{d}$ must be larger than an “inaccessibly small” fraction of the Hilbert space for us to be able to guarantee that it is not composed entirely or predominantly of states that do not thermalize. In this case, a rather convenient fact is that the ratio $d/D$ corresponding to a given energy shell is also measurable — e.g., it can be obtained by a suitable integral of $L_{\mathcal{H}}(t)$ [Eq. (112)], which is the Fourier transform of the probability density of energy levels.
@@
which shows that energy-band thermalization implies a decay of echoes to precisely the range of magnitudes necessary to establish energy-band thermalization in turn, indicating a strong interdependence [^16] For formal reasons having to do with the fact that in practice, $W,V\to 1$ but $w_{0},v_{0}\to 0$, we cannot choose the same averaging duration and the corresponding energy window for both Theorems, which would show the exact equivalence of energy-band thermalization and decaying echoes. Instead, their interdependence is to be interpreted as follows: energy-band thermalization over a certain energy scale implies decaying echoes over a longer timescale (compared to the inverse energy scale), which in turn implies energy-band thermalization with a narrower energy scale and so on. Intuitively speaking, this is “almost” an equivalence, which we expect can be made rigorous in some limit of infinite time averages after first taking a thermodynamic or classical limit, analogous to equivalence statements between autocorrelators and global eigenstate thermalization in semiclassical wavefunction ergodicity theorems [Zelditch, Sunada]..
@@
Another aspect of thermalization such a strategy can account for is the presence of conserved quantities. In addition to energy, the thermal value may also depend on some conserved charge(s) $\hat{Q}$ that commutes with the Hamiltonian, $[\hat{H},\hat{Q}]=0$ (for simplicity, we will illustrate our arguments with a single conserved charge). In that case, we will have to restrict our subspace not just to e.g. an energy interval, but to an energy shell within this interval with the conserved charge in a range $\mathcal{Q}$ centered at $Q_{c}$ with width $\Delta Q$. Here, in addition to evolving the state with the Hamiltonian $\hat{H}$, one must also apply the symmetry transformation generated by the conserved charge, through echoes such as
@@ 7.1 Cloned Hilbert spaces
So far, we have only considered time-averaged thermalization, with the promise of generalizing these results to thermal equilibrium. In this section, we will develop this generalization by showing that an observable attaining thermal equilibrium for almost all times in a given state is equivalent to the thermalization o.a. of a “cloned” observable in a “cloned state” in two copies of the Hilbert space. This is a quantum analogue of a classical result on weak mixing of product systems [SinaiCornfeld] — where a dynamical system is weak-mixing if and only if its product with itself (a doubled system) is ergodic.
@@
for some small $\epsilon$. In the $D^{2}$-dimensional doubled Hilbert space $\mathcal{H}\otimes\mathcal{H}$, we define the following “cloned” operators [^17] We use the terminology “cloned” instead of mere doubling, because we specifically want to restrict these considerations to operators or states of the form $\hat{A}\otimes\hat{A}$ or $\hat{\rho}\otimes\hat{\rho}$, the latter corresponding to cloning the quantum state [NielsenChuang]. In contrast, a mere doubled Hilbert space also admits e.g. operators $\hat{A}\otimes\hat{B}$ that are not of the “cloned” form.:
@@
$\widehat{\delta A}_{2}$ $\equiv\left(\hat{A}-A_{\mathrm{th}}\hat{\mathds{1}}\right)\otimes\left(\hat{A}-A_{\mathrm{th}}\hat{\mathds{1}}\right),$ (127)
@@
Therefore, the condition for the thermal equilibrium of $\hat{A}$ in the state $\hat{\rho}$ is that $\widehat{\delta A}_{2}$ thermalizes o.a. to 00 in the state $\hat{\rho}_{2}$:
@@ 7.2 Cloned energy-band thermalization $\implies$ thermal equilibrium
For example, using Definition 4.1 for energy-band thermalization in the subspace $\Sigma_{d}\otimes\Sigma_{d}$ of $\mathcal{H}\otimes\mathcal{H}$ gives the criterion [^18] This differs from what is called “quantum mixing” in the semiclassical chaos literature [ZelditchMixing, Zelditch], associated with a weakly mixing classical system [HalmosErgodic, SinaiCornfeld, Sinai1976], which corresponds to shifting a shrinking energy band with $\Delta E\to 0$ to be centered around some $\delta E=\delta E_{c}$ as opposed to $\delta E=0$. We prefer to work with cloned energy-band thermalization for general quantum systems due to its more direct implications for thermal equilibrium over a discrete set of times, while it is not clear to us if the aforementioned semiclassical notion will generalize to be similarly accessible except with a continuum of times.:
@@
In particular, let us take advantage of the fact that $w(t)\geq 0$ (by assumption), and the integrand on the left hand side of Eq. (136) is nonnegative as well. Then, for some $10<\kappa<$ 1 (with the expectation that $\kappa<1$ is small, but not vanishingly small), the set of times $W_{\text{ex}}[\kappa]\subseteq\mathbb{R}$, in which
@@
Then, for any constant $10<\kappa<$ 1, there exists a (possibly empty) set of times $W_{\text{ex}}[\kappa]$ spanning at most a small weighted fraction of available times [Eq. (138)]:
@@ 7.3 Cloned energy-band thermalization from quantum dynamical echoes
To see why this should work, we can write the expectation value of the cloned version of $(\hat{\Pi}_{A}-\langle\hat{\Pi}_{A}\rangle_{\Sigma_{d}}\hat{\mathds{1}})$ in $\hat{\Gamma}_{A}^{\text{shell}}$ in the energy eigenbasis :
$\operatorname{Tr}_{\mathcal{H}\otimes\mathcal{H}}$ $\left[\hat{\Gamma}_{A}^{\text{shell}}\left(\hat{\Pi}_{A}-\langle\hat{\Pi}_{A}\rangle_{\Sigma_{d}}\hat{\mathds{1}}\right)\otimes\left(\hat{\Pi}_{A}-\langle\hat{\Pi}_{A}\rangle_{\Sigma_{d}}\hat{\mathds{1}}\right)\right]$
@@
$\left\lvert\langle E_{m}\rvert\left(\hat{\Pi}_{A}-\langle\hat{\Pi}_{A}\rangle_{\Sigma_{d}}\hat{\mathds{1}}\right)\lvert E_{n}\rangle\right\rvert^{2}\left\lvert\langle E_{k}\rvert\left(\hat{\Pi}_{A}-\langle\hat{\Pi}_{A}\rangle_{\Sigma_{d}}\hat{\mathds{1}}\right)\lvert E_{\ell}\rangle\right\rvert^{2}\right\}.$ (143)
@@
$\operatorname{Tr}_{\mathcal{H}\otimes\mathcal{H}}$ $\left[\hat{\Gamma}_{A}^{\text{shell}}\left(\hat{\Pi}_{A}-\langle\hat{\Pi}_{A}\rangle_{\Sigma_{d}}\hat{\mathds{1}}\right)\otimes\left(\hat{\Pi}_{A}-\langle\hat{\Pi}_{A}\rangle_{\Sigma_{d}}\hat{\mathds{1}}\right)\right]$
@@ 8 Discussion: A sketch of experimental measurement protocols
Our interest in this section is to demonstrate the theoretical feasibility of such measurements with resources that scale at most linearly in the system size $N_{S}$ (for few-body observables). We do not perform a detailed analysis of additional $O(1)$ overheads and classical computation costs that may be incurred by the time averaging procedure, and leave a concrete experimental proposal in specific platforms with a rigorous, quantitative analysis of the measurement budget for future work. A schematic of the protocol discussed here is illustrated in Fig. 6.
@@
For example, given the computational basis states |$\lvert\{s_{j}\}_{j=1}^{N}\rangle$ of the $N$ qubits (where $s_{j}\in\{0,1\}$), $\lvert a\rangle_{S}$ could be any computational basis state of the chosen $N_{S}$-qubit subsystem, with the state of the remaining qubits being ignored. In this case, $\hat{\Pi}_{A}$ may be measured via projective measurements in the computational basis of $\mathcal{H}_{S}$, while not performing any measurements on $\mathcal{H}_{E}$.
@@
A simple strategy to achieve the maximally mixed state [NielsenChuang] is to maximally entangle $\mathcal{H}_{E}$ with a duplicate subsystem $\mathcal{H}_{R}$, also of (at least) $N_{E}$ qubits [^19] In this case, the setting is similar to one of relevance to the fast scrambling problem [HaydenPreskill].. This can be done by maximally entangling each qubit in $\mathcal{H}_{E}$ (say, indexed by $j$) with a corresponding qubit in $\mathcal{H}_{R}$, thereby generating an EPR pair [NielsenChuang] of these qubits; as entangling any pair of qubits requires only $O(1)$ steps (e.g., a Hadamard gate followed by a CNOT gate on each pair initialized to the $\lvert 00\rangle$ state), we expect the preparation of the maximally mixed state $\hat{\mathds{1}}_{E}/D_{E}$ to require only $O(N_{E})$ gates. In fact, we expect this step to be the most expensive system-independent step in our protocol in terms of the scaling of gate complexity with system size (except for implementing the actual dynamics, which depends on the Hamiltonian $\hat{H}$). As all of these maximally entangled pairs can be generated in parallel, the time complexity of this step remains $O(1)$.
@@
Now, let us consider the measurement of echoes. The echo involves the trace of a quantum mechanical operator, which is not as straightforward to measure as a return probability. We are aware of two measurement frameworks, which can be implemented in present-day experimental platforms, that can measure the squared magnitude of a trace — specifically, the spectral form factor $K(t)=\lvert L_{\mathcal{H}}(t)\rvert^{2}$ [see also Eq. (113)]. The first is by implementing controlled dynamics with an auxiliary qubit [SFFmeas], and the second by using randomized measurements [pSFF]. As the former is conceptually simpler, and also related to phase measurement algorithms [NielsenChuang], we will describe measurements of echoes using an auxiliary qubit, generalizing the discussion in Ref. [SFFmeas]. It may also be possible to adapt other measurement strategies free of auxiliary qubits [KastnerTwoTimeAmplitudes], provided there are more intrinsic ways to identify a preferred “zero” of energy $E_{c}=0$ in the system.
@@
    $\hat{C}_{U}=\hat{\Pi}_{A}\otimes\hat{\mathds{1}}_{\text{aux}}\left(e^{-i\hat{H}t_{1}}\otimes\lvert 0\rangle_{\text{aux}}\langle 0\rvert+e^{-i\hat{H}t_{2}}\otimes\lvert 1\rangle_{\text{aux}}\langle 1\rvert\right),$ (161)
    which only requires the implementation of controlled Hamiltonian dynamics (which implicitly sets a 00 value of the energy $E$ as the one that causes no phase changes in the auxiliary qubit, relative to which we must choose the center $E_{c}$ of the energy shell of interest), with the measurement of $\hat{\Pi}_{A}$ being performed just on $\mathcal{H}$ subsequent to the evolution. One way to think of such a controlled operation with different times is, e.g., if $t_{2}>t_{1}>0$, then one could implement the unitary $\exp[-i\hat{H}t_{1}]$ on $\mathcal{H}$ without any control, then apply a controlled unitary $\exp[-i\hat{H}(t_{2}-t_{1})]\otimes\lvert 1\rangle_{\text{aux}}\langle 1\rvert$ only for the remaining duration $(t_{2}-t_{1})$. Additionally, if reverse time evolution is not feasible for a choice of $t_{1}<0$ or $t_{2}<0$, we may move the corresponding dynamics to the other term to ensure that strictly positive evolution; for example, we may also set (if $t_{2}<0$)
@@
    This is essentially the proposal for spectral form factors in Ref. [SFFmeas], where $\lvert L_{\mathcal{H}}(t)\rvert^{2}$ requires a measurement of $\langle\sigma_{\text{aux}}^{x}\rangle^{2}+\langle\sigma_{\text{aux}}^{y}\rangle^{2}$, but the spectral function $L_{\mathcal{H}}(t)$ itself is obtained by measuring $\langle\sigma_{\text{aux}}^{+}\rangle$ via Eq. (158).
@@
We emphasize that except for the implementation of controlled dynamics, there is no specific aspect for these measurements that intrinsically scale with the system size, except for state preparation. The number of measurements required will be determined by the desired accuracy $\epsilon_{A}$ of the echoes, which we expect to scale with $N_{S}$ in most practical situations (assumed here to be $O(1)$) rather than $N_{E}$. For an example with a concrete proposal for implementing controlled Hamiltonian dynamics, specifically in Rydberg atoms with an auxiliary atomic clock qubit, see Ref. [SFFmeas].
@@
$\widetilde{v}\left(E+\frac{2\pi\ell}{\delta t}-E_{c}\right)=\widetilde{v}(E-E_{c}),\;\;\ \text{ for }\ell\in\mathbb{Z}.$ (167)

### v2

@@ 5.1 Completely positive time averages
    Through a relation between non-negativity conditions and Fourier transforms [72, 71], this implies:
@@
    $\mathbbR.\widetilde{w}_{+}(\delta E)\geq 0,\text{ for all }\ \delta E\in\mathbb{R}$. (83)
@@
where $\overline{\Theta}(x)=1$ if $x$ is true and 0 otherwise. However, the conventional time average $w(t)=(2T)^{-1}\overline{\Theta}(\lvert t\rvert\leq T)$ is not completely positive. Another important class of functions that are completely positive, which we will not discuss until Sec. 8, are discrete-time versions of Eq. (84) which are more accessible in an experimental setting.
Additionally, for some $0<W<1$, let $\Delta E_{W}>0$ be such that:
@@ 5.2 Energy-band thermalization from autocorrelators
Now, we will explicitly show that energy-band thermalization can be inferred from autocorrelators with a completely positive weighting function $w_{+}(t)$ which could, in principle, have support only on a finite range (or even finite set) of times. The intuition behind this relation has been described following Eq. (78), and we begin here by filling in the technical details. Formally, we will adapt a proof strategy that appeared in Refs. [40, 41, 46], generalize it to time averages with weights $w_{+}(t)$ and energy-band thermalization in place of (weak) eigenstate thermalization, and excise from it all references to the classical limit or limits such as $T\to\infty$ (or $\Delta E\to 0$). The following proposition results:
@@ 5.3 Bypassing energy levels: Global thermalization from autocorrelators
We emphasize, however, that it may still be possible for a complete basis of states that is not an eigenbasis of $\hat{\Pi}_{A}$ to thermalize this observable without implying the thermalization of the autocorrelator, nor therefore the thermalization of all other states. This strongly implicates $\hat{\rho}_{A}$ or any of its eigenbases as the most direct determiners of thermalization — which is a more accessible set of states than the complete set of all product states implicated in a slightly different setting [58] — while the observable may continue to thermalize in certain other bases of states without any implications for global thermalization [16, 74].
@@ 6.1 Motivation: general structure of accessible thermalization
[…]
@@
[…]
@@
[…]
@@ 6.2 Echo dynamics for accessing energy shells
As noted in Sec. 1, $L_{AA}(t_{1},t_{2})$ is a coarse-grained variant of the Loschmidt echo, which measures quantum interference effects between two possible evolution times of the state $\hat{\rho}_{A}$ as witnessed by the observable $\hat{\Pi}_{A}$. $L_{A}(t)$ is a stranger quantity that acts on the initial state $\hat{\rho}_{A}$ but has no observable, and we would like to refer to it as a “one-sided” echo; alternatively, it can be interpreted as an echo of the maximally mixed state with the observable $\hat{\Pi}_{A}$. Finally, $L_{\mathcal{H}}(t)$ is the spectral function of the Hamiltonian $\mathcal{H}$, measuring the Fourier transform of its probability density of energy levels (normalized to $1$); its squared magnitude is the spectral form factor [62, 75]:
@@
Finally [^15] Here, we assume that $\mathcal{E}_{V}$ will typically be a closed set, while we have taken the energy-band to be an open set/interval; this is purely a choice of convention., given some $V$ such that $0<V<1$ let us introduce $\mathcal{E}_{V}$ as a set of energy values containing $E_{c}$ in which $\widetilde{v}_{+}(E-E_{c})$ is as large as $V$:
@@ 6.3 Energy shell thermalization from quantum dynamical echoes
in a system of $N$ particles, for some $\alpha>0$, for accessibility purposes. Intuitively, this is in line with what we saw for global thermalization in Sec. 5: a vanishingly small fraction of states may fail to thermalize with accessible resources (Sec. 4) even with autocorrelator thermalization, and to have higher resolution and restrict this fraction (perhaps all the way to <1/D<1/D) one must allow “inaccessible” values of $\epsilon_{A}$. Here, we see that the energy shell $\Sigma_{d}$ must be larger than an “inaccessibly small” fraction of the Hilbert space for us to be able to guarantee that it is not composed entirely or predominantly of states that do not thermalize. In this case, a rather convenient fact is that the ratio $d/D$ corresponding to a given energy shell is also measurable — e.g., it can be obtained by a suitable integral of $L_{\mathcal{H}}(t)$ [Eq. (112)], which is the Fourier transform of the probability density of energy levels.
@@
which shows that energy-band thermalization implies a decay of echoes to precisely the range of magnitudes necessary to establish energy-band thermalization in turn, indicating a strong interdependence [^16] For formal reasons having to do with the fact that in practice, $W,V\to 1$ but $w_{0},v_{0}\to 0$, we cannot choose the same averaging duration and the corresponding energy window for both Theorems, which would show the exact equivalence of energy-band thermalization and decaying echoes. Instead, their interdependence is to be interpreted as follows: energy-band thermalization over a certain energy scale implies decaying echoes over a longer timescale (compared to the inverse energy scale), which in turn implies energy-band thermalization with a narrower energy scale and so on. Intuitively speaking, this is “almost” an equivalence, which we expect can be made rigorous in some limit of infinite time averages after first taking a thermodynamic or classical limit, analogous to equivalence statements between autocorrelators and global eigenstate thermalization in semiclassical wavefunction ergodicity theorems [46, 41]..
@@
Another aspect of thermalization such a strategy can account for is the presence of conserved quantities. In addition to energy, the thermal value may also depend on some conserved charge(s) $\hat{Q}$ that commutes with the Hamiltonian, $\hat{H},\hat{Q}]=0$ (for simplicity, we will illustrate our arguments with a single conserved charge). In that case, we will have to restrict our subspace not just to e.g. an energy interval, but to an energy shell within this interval with the conserved charge in a range $\mathcal{Q}$ centered at $Q_{c}$ with width $\Delta Q$. Here, in addition to evolving the state with the Hamiltonian $\hat{H}$, one must also apply the symmetry transformation generated by the conserved charge, through echoes such as
@@ 7.1 Cloned Hilbert spaces
So far, we have only considered time-averaged thermalization, with the promise of generalizing these results to thermal equilibrium. In this section, we will develop this generalization by showing that an observable attaining thermal equilibrium for almost all times in a given state is equivalent to the thermalization o.a. of a “cloned” observable in a “cloned state” in two copies of the Hilbert space. This is a quantum analogue of a classical result on weak mixing of product systems [6] — where a dynamical system is weak-mixing if and only if its product with itself (a doubled system) is ergodic.
@@
for some small $\epsilon$. In the $D^{2}$-dimensional doubled Hilbert space $\mathcal{H}\otimes\mathcal{H}$, we define the following “cloned” operators [^17] We use the terminology “cloned” instead of mere doubling, because we specifically want to restrict these considerations to operators or states of the form $\hat{A}\otimes\hat{A}$ or $\hat{\rho}\otimes\hat{\rho}$, the latter corresponding to cloning the quantum state [52]. In contrast, a mere doubled Hilbert space also admits e.g. operators $\hat{A}\otimes\hat{B}$ that are not of the “cloned” form.:
@@
[…]
@@
Therefore, the condition for the thermal equilibrium of $\hat{A}$ in the state $\hat{\rho}$ is that $\widehat{\delta A}_{2}$ thermalizes o.a. to 0 in the state $\hat{\rho}_{2}$:
@@ 7.2 Cloned energy-band thermalization $\implies$ thermal equilibrium
For example, using Definition 4.1 for energy-band thermalization in the subspace $\Sigma_{d}\otimes\Sigma_{d}$ of $\mathcal{H}\otimes\mathcal{H}$ gives the criterion [^18] This differs from what is called “quantum mixing” in the semiclassical chaos literature [45, 46], associated with a weakly mixing classical system [5, 6, 7], which corresponds to shifting a shrinking energy band with $\Delta E\to 0$ to be centered around some $\delta E=\delta E_{c}$ as opposed to $\delta E=0$. We prefer to work with cloned energy-band thermalization for general quantum systems due to its more direct implications for thermal equilibrium over a discrete set of times, while it is not clear to us if the aforementioned semiclassical notion will generalize to be similarly accessible except with a continuum of times.:
@@
[…]
@@
Then, for any constant $0<\kappa<1$, there exists a (possibly empty) set of times $W_{\text{ex}}[\kappa]$ spanning at most a small weighted fraction of available times [Eq. (138)]:
@@ 7.3 Cloned energy-band thermalization from quantum dynamical echoes
[…]
@@
[…]
@@
[…]
@@ 8 Discussion: A sketch of experimental measurement protocols
Our interest in this section is to demonstrate the theoretical feasibility of such measurements with resources that scale at most linearly in the system size $N$ (for few-body observables). We do not perform a detailed analysis of additional $O(1)$ overheads and classical computation costs that may be incurred by the time averaging procedure, and leave a concrete experimental proposal in specific platforms with a rigorous, quantitative analysis of the measurement budget for future work. A schematic of the protocol discussed here is illustrated in Fig. 6.
@@
For example, given the computational basis states $\lvert\{s_{j}\}_{j=1}^{N}\rangle$ of the $N$ qubits (where $s_{j}\in\{0,1\}$), $\lvert a\rangle_{S}$ could be any computational basis state of the chosen $N_{S}$-qubit subsystem, with the state of the remaining qubits being ignored. In this case, $\hat{\Pi}_{A}$ may be measured via projective measurements in the computational basis of $\mathcal{H}_{S}$, while not performing any measurements on $\mathcal{H}_{E}$.
@@
A simple strategy to achieve the maximally mixed state [52] is to maximally entangle $\mathcal{H}_{E}$ with a duplicate subsystem $\mathcal{H}_{R}$, also of (at least) $N_{E}$ qubits [^19] In this case, the setting is similar to one of relevance to the fast scrambling problem [76].. This can be done e.g. by maximally entangling each qubit in $\mathcal{H}_{E}$ (say, indexed by $j$) with a corresponding qubit in $\mathcal{H}_{R}$, thereby generating an EPR pair [52] of these qubits; as entangling any pair of qubits requires only $O(1)$ steps (e.g., a Hadamard gate followed by a CNOT gate on each pair initialized to the $\lvert 00\rangle$ state), we expect the preparation of the maximally mixed state $\hat{\mathds{1}}_{E}/D_{E}$ to require only $O(N_{E})$ gates. In fact, we expect this step to be the most expensive system-independent step in our protocol in terms of the scaling of gate complexity with system size (except for implementing the actual dynamics, which depends on the Hamiltonian $\hat{H}$). As all of these maximally entangled pairs can be generated in parallel, the time complexity of this step remains $O(1)$.
@@
Now, let us consider the measurement of echoes. The echo involves the trace of a quantum mechanical operator, which is not as straightforward to measure as a return probability. We are aware of two measurement frameworks, which can be implemented in present-day experimental platforms, that can measure the squared magnitude of a trace — specifically, the spectral form factor $K(t)=\lvert L_{\mathcal{H}}(t)\rvert^{2}$ [see also Eq. (113)]. The first is by implementing controlled dynamics with an auxiliary qubit [22], and the second by using randomized measurements [23]. As the former is conceptually simpler, and also related to phase measurement algorithms [52], we will describe measurements of echoes using an auxiliary qubit, generalizing the discussion in Ref. [22]. It may also be possible to adapt other measurement strategies free of auxiliary qubits [77], provided there are more intrinsic ways to identify a preferred “zero” of energy $E_{c}=0$ in the system.
@@
[…]
    which only requires the implementation of controlled Hamiltonian dynamics (which implicitly sets a 0 value of the energy $E$ as the one that causes no phase changes in the auxiliary qubit, relative to which we must choose the center $E_{c}$ of the energy shell of interest), with the measurement of $\hat{\Pi}_{A}$ being performed just on $\mathcal{H}$ subsequent to the evolution. One way to think of such a controlled operation with different times is, e.g., if $t_{2}>t_{1}>0$, then one could implement the unitary $\exp[-i\hat{H}t_{1}]$ on $\mathcal{H}$ without any control, then apply a controlled unitary $\exp[-i\hat{H}(t_{2}-t_{1})]\otimes\lvert 1\rangle_{\text{aux}}\langle 1\rvert$ only for the remaining duration $(t_{2}-t_{1})$. Additionally, if reverse time evolution is not feasible for a choice of $t_{1}<0$ or $t_{2}<0$, we may move the corresponding dynamics to the other term to ensure that strictly positive evolution; for example, we may also set (if $t_{2}<0$)
@@
    This is essentially the proposal for spectral form factors in Ref. [22], where $\lvert L_{\mathcal{H}}(t)\rvert^{2}$ requires a measurement of $\langle\sigma_{\text{aux}}^{x}\rangle^{2}+\langle\sigma_{\text{aux}}^{y}\rangle^{2}$, but the spectral function $L_{\mathcal{H}}(t)$ itself is obtained by measuring $\langle\sigma_{\text{aux}}^{+}\rangle$ via Eq. (158).
@@
We emphasize that except for the implementation of controlled dynamics, there is no specific aspect of these measurements that intrinsically scale with the system size, except for state preparation. The number of measurements required will be determined by the desired accuracy $\epsilon_{A}$ of the echoes, which we expect to scale with $N_{S}$ in most practical situations (assumed here to be $O(1)$) rather than $N_{E}$. For an example with a concrete proposal for implementing controlled Hamiltonian dynamics, specifically in Rydberg atoms with an auxiliary atomic clock qubit, see Ref. [22].
Additionally, in implementations where the qubits in $\mathcal{H}$ are each part of a larger local Hilbert space (e.g., $2$-level subspaces of $4$-level systems), it is possible to directly “upgrade” any isolated Hamiltonian evolution $e^{-i\hat{H}t}$ of the qubits to a variant $\hat{C}_{U}$ controlled by an external qubit, without any explicit knowledge of $\hat{H}$ or interfering with the isolated dynamics of qubits. This may be done by rapidly applying simpler controlled operations that swap each local qubit subspace with an orthogonal local subspace that is not subject to any time evolution, before and after the Hamiltonian dynamics of qubits occurs for the desired length of time over which it must be controlled [78]. This may further allow the implementation of such echo measurements in experimental platforms where controlled variants of a given Hamiltonian $\hat{H}$ cannot be “natively” implemented.
@@
$\mathbbZ.\widetilde{v}\left(E+\frac{2\pi\ell}{\delta t}-E_{c}\right)=\widetilde{v}(E-E_{c}),\;\;\ \text{ for }\ell\in\mathbb{Z}$. (167)